Given: π‘Œ = 𝐢+𝐼,𝐢 = 50+0.8π‘Œπ‘‘,π‘Žπ‘›π‘‘ 𝐼0 = 50.

(a) Graph the consumption function.

(b) Graph the aggregate demand function, 𝐢+𝐼0.

(c) Find the equilibrium level of income from the graph.


1
Expert's answer
2021-05-31T09:10:30-0400

(a) The consumption function, or Keynesian consumption function, is an economic formula that represents the functional relationship between total consumption and gross national income.

The graph of this function is a line, and it will start from the point (0; 50) crossing the C-axis (y-axis) and will cross the 45-degree line.

(b) The aggregate demand function is a downward-sloping line in the graph with real GDP at x-axis and price level at y-axis.

(c) The equilibrium level of income is Y = C + I = 50 + 0.8Yd + 50 = 100 + 0.8Yd.

Yd = Y - T, so:

Y = 100 + 0.8(Y - T) = 100 + 0.8Y - 0.8T,

0.2Y = 100 - 0.8T,

Y = 500 - 4T, where T is taxes.